Hi,I need some help solving this equation please
The ordered pair (3,2) is a solution of the equation 2x-5y=?

Answer by stanbon(75887)

Hi,I need some help solving this equation please
The ordered pair (3,2) is a solution of the equation 2x-5y=?
-----------------------------
Since 2x-5y = ?,
---
? = 2*3 - 5*2
? = 6-10
? = -4
=============
Cheers,
Stan H.